# Foundational GPT Model for MEG

### v1

@@ A.3 Generated covariance
On most metrics presented in the main paper FlatGPT2 performs considerably worse than ChannelGPT2. As the PSD is a channel-independent measure, we also looked at generated data covariance which captures the interactions between different channels (Figure 19). This reveals that the only model capable of closely matching the data covariance is FlatGPT2. All other models produce data with covariances much closer to 0. This is perhaps expected for channel-independent models which generate data independently for each channel, but somewhat surprising for WavenetFullChannelMix. Even though FlatGPT2 may not produce accurate spectral data, by having information about other channels in the input it does an excellent job at capturing covariance. This highlights the trade-offs between different model architectures.
A.4 FlatGPT2 on group data
Unfortunately, even scaling FlatGPT2 did not improve evoked generation. However, we did find that the spectral content of the generated data matched the real data much better than the single-subject version of FlatGPT2 (Figure 20). FlatGPT2-group seemed to scale particularly well with model size as larger models achieved lower and lower loss, improving test accuracy by multiple folds (16.1% top-1 and 40.1% top-5 accuracy) over single-subject FlatGPT2 (3% top-1 and 10.8% top-5 accuracy). This is interesting behaviour compared to ChannelGPT2-group which did not improve much on our forecasting metrics. It remains to be seen whether even more data and larger models are needed to make this type of architecture viable.
@@ A.4 FlatGPT2 on group data
(b) FlatGPT2
(c) FlatGPT2-group
Figure 20: Comparison of generated data PSD across data single-subject FlatGPT2 and FlatGPT2-group. Each line represents a different MEG channel.

### v2

@@ A.10 Generated covariance
As the PSD is a channel-independent measure, we also looked at generated data covariance which captures the interactions between different channels (Figure 17). All models produce data with covariances much closer to 0 than real data. This is perhaps expected for channel-independent models which generate data independently for each channel, but somewhat surprising for WavenetFullChannelMix.
A.11 GPT2MEG-group evoked analysis
(a) Visual channel (MEG2332)
(b) Frontal channel (MEG0111)
Figure 18: Comparison of evoked responses averaged across all subjects in the real MEG data (blue) and across all subjects in the generated data from the GPT2MEG-group model trained on all subjects (orange). The horizontal axis encompasses 1 second, stimulus onset is at 0 seconds and stimulus offset is at 0.5 seconds. Shading indicates 95% confidence interval of the trial mean. ChannelGPT refers to GPT2MEG.
To test our hypothesis regarding GPT2MEG-group generating more of an average across subjects, we generated data for all subjects (using appropriate subject embeddings) and compared the grand average evoked responses with those extracted from the MEG data of all subjects. Two channels are plotted in Figure 18. The evoked response averaged over all subjects is much noisier because of the high between-subject variability. However, we can see that indeed GPT2MEG-group can generate this well, perhaps slightly smoother than the real data. Comparing these plots with Figure 5, it is also clear that it adapts its generation well to a specific subject compared to the group average. Further evoked activity comparisons based on HMM state timecourses are shown below.
Finally, we examine the variability in state time courses over individual trials. For this we inferred an 8-state HMM on the real data of a single subject. Then, using the HMM observation models held fixed from the real data, we also estimated the state timecourses on both the single-subject GPT2MEG and GPT2MEG-group generated data (with the appropriate subject embedding). Holding the observation models fixed allows us to obtain matched states. We hypothesised that even if the average evoked responses are similar to the real data, GPT2 may not be able to generate trials with variability in the temporal activation of states. Figure 19 shows that this is indeed true for the single-subject GPT2MEG generated data. GPT2MEG-group responses seem to include much higher temporal variability in state activations, though still falling short of the real data. This indicates that the model can capture some trial-to-trial variability through its exposure to multiple subjects, but has difficulty fully matching the complexity of real neural data. More data may be needed to improve this aspect of generation.
@@ A.11 GPT2MEG-group evoked analysis
(b) GPT2MEG
(c) GPT2MEG-group
Figure 19: Comparison of the trial-level variability in the evoked HMM state timecourses from an HMM inferred on (a) real MEG data, and on data generated from the (b) GPT2MEG model trained on a single sample subject and (c) GPT2MEG-group trained on all subjects. Different colours represent different states (matched across models). Individual trials however are not matched and we cannot compare the plots at the trial-level, only as an aggregate visualisation of variability across trials.
To further test alignment between group-level evoked responses from a multi-channel perspective, we: 1) inferred an HMM on the real MEG data across all subjects, 2) generated data across all subjects from the trained GPT2MEG-group model, then 3) estimated HMM state timecourses on the generated data, but using the observation models from the HMM inferred in step 1. By holding the observation models fixed we can directly match the evoked HMM state timecourses between the real MEG and generated timeseries. We trained an amplitude-envelope HMM (AE-HMM) with 6 states (Quinn et al., 2019)) and show results in Figure 20. Two states that show strong activation during real task data show similar temporal signatures and amplitude changes in the generated data, albeit slightly noisier. In the generated data, there are two additional states which seem to get activated during the trial. This indicates that while GPT2MEG-group can capture some of the state-level dynamics, there is room for improvement.
(a) Data
(b) GPT2MEG-group
Figure 20: Comparison of the evoked HMM state timecourses (averaged across all subjects); from (a) an HMM inferred on real MEG data, and from (b) data generated from the GPT2MEG-group model trained on all subjects, but where the observation models from the HMM inferred on the real MEG was used. Note that state indices are matched between the two plots, as the same fitted HMM model was used.
A.12 GPT2MEG adapts generation to different trial lengths
We evaluated the model’s ability to adapt to different trial durations. The results reported thus far are for a GPT2MEG trained on trials lasting 0.5 seconds. We generated data using the same fitted GPT2MEG model but with trial durations of 0.2 s and 0.8 s. As shown in Figure 21, GPT2MEG accurately adapted to the shorter and longer trials. The evoked responses matched the expected time-courses, with appropriate truncation or lack of second peaks due to stimulus offset. This demonstrates the model’s ability to generalise to varied trial durations despite being trained on a fixed duration.
Figure 21: Evoked responses generated by GPT2MEG for a single sample subject, for trials of 0.2 s (orange), 0.5 s (blue), and 0.8 s (green). While the model was trained only on data containing trials of 0.5 s, it shows an ability to adapt appropriately to the different durations. The plotted channel is in the visual area (MEG2332). ChannelGPT refers to GPT2MEG.